# Music Foundation Model as Generic Booster for Music Downstream Tasks

### v2

@@ 1 Introduction
A foundation model is a pre-trained model developed on a large-scale dataset that can be adapted for a variety of downstream tasks (Bommasani et al., 2021). Several language processing models (Radford et al., 2021; Brown et al., 2020; Devlin et al., 2018; Team et al., 2024) are considered foundation models due to their ability to unify all language tasks as sequence prediction tasks, effectively addressing multiple tasks with a single model. These foundation models have gained significant attraction and are widely used in everyday applications. In contrast, a powerful music foundation model capable of handling various music downstream tasks for music production is lacking. We categorize the tasks that a music foundation model primarily addresses into two types: understanding tasks, such as tagging and transcription, and generative tasks, such as mixing and mastering.
@@
We extended the methodology from MIR tasks to generic music downstream tasks and hypothesize that the representation structure of foundation models is crucial in this case. Specifically, we propose that hierarchical representations, which divide information of varying granularity into different levels of embedding, are expected to provide efficient information hierarchy for all downstream tasks including both understanding and generative tasks. We empirically verify this hypothesis in Section 4. In contrast, music foundation models that have been applied to boost music downstream tasks do not have such a hierarchical structure. For example, Jukebox (Castellon et al., 2021; Donahue et al., 2022) is trained to have multi-level representation inspired from hierarchical latent representation (Razavi et al., 2019); however, each level is independently trained. RQ-VAEs (Yang et al., 2023; Li et al., 2023b) learn factorized representation that has a self-organized coarse-to-fine structure, however, they are not hierarchical.
In accordance with the aforementioned hypothesis, we outline this study as follows. We propose and train our music foundation model, SoniDo (meaning sound in Spanish), on a high-fidelity internal dataset***The rights of this internal dataset are trained on licensed content only. Except for as specifically authorized by the rights owner, the rights owner expressly prohibits and has opted out of any text or data mining, web scraping or similar, reproductions, extractions or uses, of its content for any purposes, including in relation to training, developing, or commercializing any Al System. to establish a task-agnostic feature extraction pipeline. SoniDo is a generative model consisting of a multi-level transformer with a multi-level hierarchical encoder. With proper pre-processing, we infuse its intermediate representation as features to task-specific models on various music downstream tasks with data augmentation. Moreover, for understanding tasks, we proposed an on-the-fly data augmentation method called token-out to avoid the overfitting problem. We conducted a performance evaluation by benchmarking with representative tasks from understanding to generative tasks: music tagging, music transcription, music source separation, and music mixing.
Table 1: Performance overview of applying extracted features to various music downstream tasks. Bold: best, underline: second best. Results obtained with different evaluation protocols are marked with asterisks.
@@
CALM w/
Jukebox-5B
MERT Task-Specific SOTA
@@
Music Transcription MAPS Frame F1 83.92 82.94 81.53 - - 82.89 (Toyama et al., 2023)
Note F1 86.45 85.97 85.14 - - 85.14
Note w/ Offset F1 68.27 68.27 66.28 - - 66.34
@@
51.34 50.42 48.69 - - 48.20
Source Separation MUSDB18 SDR (bass) 9.50 8.86 8.17 4.9 5.6 11.31 (Lu et al., 2024)
SDR (drums) 8.65 8.03 7.50 4.1 3.6 9.49
SDR (other) 5.91 5.59 5.54 2.7 3.0 7.73
SDR (vocals) 8.07 7.57 7.66 5.1 5.3 10.66
MDXDB21 hidden SDR (bass) 8.14 7.44 7.40 - - 7.86 (Rouard et al., 2023)
SDR (drums) 8.16 8.31 7.37 - - 7.89
SDR (other) 5.21 5.26 4.93 - - 5.09
SDR (vocals) 8.04 7.81 7.73 - - 7.70
Music Mixing MDXDB21-dry hidden Stereo-Invariant 79.86 87.27 87.32 - - 82.09 (Martínez-Ramírez et al., 2022)
@@
As shown in Table 1, we test our SoniDo’s feature injection for several music downstream tasks. To the best of our knowledge, this is the first study on enhancing both understanding and generative tasks with the intermediate representation from a single model. We briefly list the major findings of this study:
@@
    We empirically show that, with a generative model that is established on hierarchical representation, its intermediate representation can serve as generic booster of various music downstream tasks.
@@ 2 Proposed Two-stage Hierarchical Model: SoniDo
To explore the effectiveness of hierarchical modeling in boosting downstream tasks, we adopt the typical two-stage generative modeling (Dhariwal et al., 2020; Copet et al., 2023; Li et al., 2023a). In the stage-1 model, we use an HQ-VAE for hierarchical representation learning that divides information into different levels on the basis of their granularity. In the stage-2 model, we use auto-regressive modeling to learn the multi-level token streams extracted from the stage-1 model. Finally, we extract features from stage-2 model, as described in Section 3.1.
@@ 2.2 Stage-2 Model: Sparse Transformers
We follow Jukebox (Dhariwal et al., 2020) to construct the networks $\bm{\Pi}$ with sparse transformers (Vaswani et al., 2017; Child et al., 2019). As illustrated in Figure 2(b), we train three autoregressive sparse transformers to model $P(\bm{Z}_{1}|\bm{y}_{\text{audio}})$, $P(\bm{Z}_{2}|\bm{Z}_{1},\bm{y}_{\text{audio}})$, and $P(\bm{Z}_{3}|\bm{Z}_{1,2},\bm{y}_{\text{audio}})$, which we refer to as top prior, middle conditional prior, and bottom conditional prior, respectively. The middle and bottom priors use the token sequences from the upper levels, with up-sampling achieved through upsampling modules, corresponding to the conditioners of Jukebox. We additionally condition each prior on $\bm{y}_{\text{audio}}$. Appendix B.2 provides further details of the stage-2 model. Appendix B.3 evaluates the common objective metrics on SoniDo.
@@ 3.2 Feature Pre-processing for Time-invariant Downstream Tasks
To prevent overfitting when using the concatenated token sequence to train the attention block. We propose an on-the-fly data augmentation method called token-out. This is inspired by SpecAugment (Park et al., 2019), PaSST (Koutini et al., 2022) and masked autoencoder (MAE) (He et al., 2022), in which a part of the input is masked before feeding into deep neural networks. The masking ratio is sampled between 0 and 100% uniformly. Our token-out augmentation follows the principles in PaSST and MAE but is applied to tokens extracted with SoniDo, as illustrated in Figure 4. As shown in Appendix C, aggregating the SoniDo features with the shallow attention layer and token-out augmentation led to performance improvement.
@@ 4 Experiments
We conducted experiments to examine the usefulness of features extracted from music foundation models on understanding and generative downstream tasks by answering the following two questions. Q1: Do extracted features have useful information for music understanding? Q2: Can extracted features boost current task-specific models for both understanding and generative tasks? To verify the generalizability of the results, we included not only the SoniDo features, but also features extracted from the two public versions of MusicGen (Copet et al., 2023), which are MusicGen Small and MusicGen Large.
To answer the first question, we selected eight music tagging tasks and music transcription as the representative for understanding tasks. We verify that the extracted features do not only contain the time-invariant information of overall musical properties but also the time-varying information of specific musical events. To answer the second question, we tested injecting extracted features to several task-specific models that covering both understanding and generative tasks. They consist of one music transcription model (Toyama et al., 2023), two source separation models (Mitsufuji et al., 2022; Fabbro et al., 2023), and two music mixing models from Martínez-Ramírez et al. (2022).
@@ 4.1 Usefulness of Extracted Features for Music Understanding
Through the following feature probing experiments, we show that useful information is contained within both the SoniDo and MusicGen features, which is consistent with previous results Castellon et al. (2021). However, for music transcription, the shallow network is still insufficient to achieve comparable results to recent SOTA models. In Section 4.2, we show that injecting extracted features on current task-specific models can boost their performance beyond SOTA.
@@ 4.1.1 Music Tagging
We test a wide range of music tagging tasks as well as the emotion regression task: MagnaTagATune (MTAT) (Law et al., 2009) for auto tagging, Nsynth (Engel et al., 2017) for pitch and instrument recognition, EmoMusic (Soleymani et al., 2013) for emotion regression, GTZAN (Tzanetakis & Cook, 2002) for genre classification, GiantSteps (Knees et al., 2015; Korzeniowski & Widmer, 2017) for musical key estimation, and VocalSet (Wilkins et al., 2018) for singer and singing technique identification. A summary of the datasets is shown in Table 9 in Appendix C. We followed the pre-processing in previous studies (Li et al., 2023b; Yuan et al., 2023) and used scikit-learn (Pedregosa et al., 2011) and mir_eval (Raffel et al., 2014) for metric computation. The average R2 of arousal and valence axis is reported for EmoMusic.
The feature pre-processing for time-invariant downstream tasks in Section 3.2 including the feature aggregation and token-out augmentation is applied for all tasks. Following common practice (Castellon et al., 2021; Li et al., 2023b), an MLP is then used to probe the aggregated features.
@@
Table 3: Music transcription. Evaluation results of feature probing using shallow back-end on MAPS (bold: best score, underline: second-best score).
@@ 4.1.2 Music Transcription
Beyond time-invariant understanding tasks, we continue the test on music transcription, which is a time-varying understanding task. Both SoniDo and MusicGen features are obtained with unconditional extraction described in Section 3.1. The dimension and time resolution of extracted features are aligned to those of the spectrogram by linear layers. Following Castellon et al. (2021), these aligned features are concatenated with the spectrogram, and a single-layer shallow back-end network is used to probe these features.
We show the transcription performance of the feature probing mentioned above on the MAPS dataset (Emiya et al., 2010) in Table 3. All the features greatly improved the note-wise F1 score compared with using the spectrogram only. This suggests that both SoniDo and MusicGen features contain useful information for time-varying understanding tasks.
@@ 4.2 Using Extracted Features to Boost Existing Task-specific Models
In Section 4.1, we showed that the extracted features contain useful knowledge for music understanding. In this section, we test both SoniDo and MusicGen features on several SOTA task-specific models, the tasks include music transcription, music source separation, and music mixing, which covered both music understanding and generative tasks. The experimental results indicate that the extracted features consistently boost the performances of task-specific models. We also observed that injecting the SoniDo features accelerated the decrement of training loss in early epochs.
@@ 4.2.1 Music Transcription: hFT-Transformer
We applied the extracted features to hFT-Transformer (Toyama et al., 2023), a SOTA music transcription model for piano on MAPS (Emiya et al., 2010), to assess whether it surpasses existing models that rely solely on the spectrogram. On the basis of the input spectrogram, hFT-Transformer estimates the frame-based note activation, along with the onset, offset, and velocity of a note (frame, onset, offset, and velocity). It is a transformer-based model consisting of two transformer encoders that work on different axes of the input and a transformer decoder in the middle of these two encoders. Following the processing pipeline in Section 4, we attempted injecting the SoniDo and MusicGen features before the 1st encoder, 2nd encoder, and decoder. We found that feature injection before the decoder yields the best result, thus we adopt this injection method in the following experiments. All the training hyperparameters were kept the same as in a previous study (Toyama et al., 2023). Further details are provided in Appendix D.3.
Table 4: Music transcription. Results of F1 scores on MAPS (bold: best score, underline: second best score). “Note” refers to note-wise estimation. First row corresponds to hFT-Transformer (Toyama et al., 2023).
@@
Following the common evaluation practice (Gardner et al., 2022; Toyama et al., 2023), we report four F1 scores: frame-wise, note-wise, note-wise with offset, and note-wise with offset and velocity using the checkpoint with the best validation F1 score. As shown in Table 4, injecting either SoniDo or MusicGen Small features improves the performance of hFT-Transformer. The performance gap is especially huge when the model is trained with a small subset of MAPS. This demonstrates the usefulness of injecting music foundation model features into downstream task models when training data are scarce. We also observed that the decrement of training loss is faster when either SoniDo or MusicGen features are injected, as shown in Figure 8 in Appendix D.3.
@@ 4.2.2 Music Source Separation: UMX, HTDemucs
We select Open-Unmix (UMX) (Stöter et al., 2019) and Demucs (HTDemucs) (Rouard et al., 2023) as the targets of feature injection. UMX estimates the time-frequency mask of the target source, which is based on recurrent neural network (RNN) blocks. HTDemucs is a hybrid model consisting of a waveform U-Net branch and spectral U-Net branch. We choose to inject the extracted features into the encoder block for UMX using a down-sampling block and in each HTDemucs branch using a cross-domain Transformer (details in Appendices E.1 and  E.3). On the basis of the observation in Section 4.2.1 and for the sake of simplicity, we inject only top-level features from SoniDo.
Table 5 lists the SDR scores on the test split of MUSDB18 Rafii et al. (2017) and the hidden split of MDXDB21 Mitsufuji et al. (2022); Fabbro et al. (2023). The details of the experiments are provided in Appendix E. Similar to the experiment discussed in Section 4.2.1, a faster loss decrement is observed, as shown in Figures 12 and 13 in Appendix E. Injecting the SoniDo features into both UMX and HTDemucs greatly boosts the separation performance for both models, even on the unseen dataset MDXDB21. It also improves the separation performance of HTDemucs when training on data corrupted by bleeding errors (SDXDB23_BleedingFabbro et al. (2023). However, the MusicGen features do not always improve the results. Injecting the MusicGen Small features improved UMX, but not for the other cases. According to the ablation study results on HTDemucs, injecting short-term Fourier transform (STFT) signal, CLAP features, or MusicGen features leads to unstable training. However, no such behavior is observed when injecting the SoniDo features into HTDemucs. We assume that performance can be improved if instability during training is avoided. As mentioned in Section 4.2.1, interpreting information contained in MusicGen Large could cost too much capacity of the downstream model and result in performance degradation.

### v3

@@ 1 Introduction
A foundation model is a pre-trained model developed on a large-scale dataset that can be adapted for a variety of downstream tasks (Bommasani et al., 2021). Several language processing models (Radford et al., 2021; Brown et al., 2020; Devlin et al., 2018; Team et al., 2024) are considered foundation models due to their ability to unify all language tasks as sequence prediction tasks, effectively addressing multiple tasks with a single model. These foundation models have gained significant attraction and are widely used in everyday applications. In contrast, a powerful music foundation model capable of handling various music downstream tasks for music production is lacking (Ma et al., 2024). We categorize the tasks that a music foundation model primarily addresses into two types: understanding tasks, such as tagging and transcription, and generative tasks, such as mixing and mastering.
@@
We extended the methodology from MIR tasks to generic music downstream tasks. To address both understanding and generative tasks, we focus on a tokenization-based generative modeling approach, which is directly analogous to foundation models in the natural language processing field. Furthermore, We hypothesize that the representation structure of foundation models is crucial in this context. Specifically, we propose that hierarchical representations, which divide information of varying granularity into different levels of embedding, are expected to provide efficient information hierarchy for all downstream tasks including both understanding and generative tasks. We empirically verify this hypothesis in Section 4. In contrast, music foundation models that have been applied to boost music downstream tasks do not have such a hierarchical structure. For example, Jukebox (Castellon et al., 2021; Donahue et al., 2022) is trained to have multi-level representation inspired from hierarchical latent representation (Razavi et al., 2019); however, each level is independently trained. RQ-VAEs (Yang et al., 2023; Li et al., 2023b) learn factorized representation that has a self-organized coarse-to-fine structure, however, they are not hierarchical.
In accordance with the aforementioned hypothesis, we outline this study as follows. We propose and train our music foundation model, SoniDo (meaning sound in Spanish), on a high-fidelity internal dataset  [^1] The rights of this internal dataset are trained on licensed content only. Except for as specifically authorized by the rights owner, the rights owner expressly prohibits and has opted out of any text or data mining, web scraping or similar, reproductions, extractions or uses, of its content for any purposes, including in relation to training, developing, or commercializing any Al System. to establish a task-agnostic feature extraction pipeline. SoniDo is a generative model consisting of a multi-level transformer with a multi-level hierarchical encoder. With proper pre-processing, we infuse its intermediate representation as features to task-specific models on various music downstream tasks with data augmentation. Moreover, for understanding tasks, we proposed an on-the-fly data augmentation called token-out to avoid overfitting. Performance evaluation was done by benchmarking with representative tasks from understanding to generative tasks: music tagging, music transcription, music source separation, and music mixing.
Table 1: Performance overview of applying extracted features to various downstream tasks. Bold: best, underline: second best. The result marked with * is obtained with a different evaluation protocol. The results marked with † are numbers reported in Castellon et al. (2021).
@@
Jukebox-5B MERT Task-Specific SOTA
@@
Music Transcription MAPS Frame F1 83.92 82.94 81.53 84.23 - 82.89 (Toyama et al., 2023)
Note F1 86.45 85.97 85.14 86.54 - 85.14
Note w/ Offset F1 68.27 68.27 66.28 68.26 - 66.34
@@
51.34 50.42 48.69 50.46 - 48.20
Source Separation MUSDB18 SDR (bass) 9.50 8.86 8.17 7.12 5.6 11.31 (Lu et al., 2024)
SDR (drums) 8.65 8.03 7.50 6.65 3.6 9.49
SDR (other) 5.91 5.59 5.54 4.77 3.0 7.73
SDR (vocals) 8.07 7.57 7.66 6.84 5.3 10.66
MDXDB21 hidden SDR (bass) 8.14 7.44 7.40 6.58 - 7.86 (Rouard et al., 2023)
SDR (drums) 8.16 8.31 7.37 6.58 - 7.89
SDR (other) 5.21 5.26 4.93 4.59 - 5.09
SDR (vocals) 8.04 7.81 7.73 7.12 - 7.70
Music Mixing MDXDB21-dry hidden Stereo-Invariant 79.86 87.27 87.32 87.97 - 82.09 (Martínez-Ramírez et al., 2022)
@@
As shown in Table 1, we test SoniDo’s feature injection for selected downstream tasks along with several baselines. To the best of our knowledge, this is the first study on enhancing both understanding and generative tasks with the intermediate representation from a single model. We briefly list our major findings:
@@
    We empirically show that, with an auto-regressive generative model that is established on hierarchical representation, its intermediate representation can serve as generic booster of various music downstream tasks.
@@ 2 Proposed Two-stage Hierarchical Model: SoniDo
To explore the effectiveness of hierarchical modeling in boosting downstream tasks, we adopt a typical two-stage generative modeling (Dhariwal et al., 2020; Copet et al., 2023; Li et al., 2023a). In stage-1, we use an HQ-VAE for hierarchical representation learning, dividing information into different levels based on granularity. In stage-2, we use auto-regressive modeling to learn the multi-level token streams extracted from the stage-1 model. Finally, features from stage-2 model are extracted as described in Section 3.1.
@@ 2.2 Stage-2 Model: Sparse Transformers
We follow Jukebox (Dhariwal et al., 2020) to construct the networks $\bm{\Pi}$ with sparse transformers (Vaswani et al., 2017; Child et al., 2019). As illustrated in Figure 2(b), we train three auto-regressive sparse transformers to model $P(\bm{Z}_{1}|\bm{y}_{\text{audio}})$, $P(\bm{Z}_{2}|\bm{Z}_{1},\bm{y}_{\text{audio}})$, and $P(\bm{Z}_{3}|\bm{Z}_{1,2},\bm{y}_{\text{audio}})$, which we refer to as top prior, middle conditional prior, and bottom conditional prior, respectively. The middle and bottom priors use the token sequences from the upper levels, with up-sampling achieved through upsampling modules, corresponding to the conditioners of Jukebox. We additionally condition each prior on $\bm{y}_{\text{audio}}$. Appendix B.2 provides further details of the stage-2 model. Appendix B.3 evaluates the common objective metrics on SoniDo.
@@ 3.2 Feature Pre-processing for Time-invariant Downstream Tasks
To prevent overfitting when using the concatenated token sequence to train the attention block, we propose an on-the-fly data augmentation method called token-out. This is inspired by SpecAugment (Park et al., 2019) and masked Transformers (Koutini et al., 2022; Zhong et al., 2023b; Comunità et al., 2024), in which a part of the input is masked before feeding into deep neural networks. Unlike prior arts, token-out is applied to the whole token sequence extracted with the multiple layers in SoniDo, as illustrated in Figure 4.The masking ratio is sampled between 0 and 100% uniformly. As shown in Appendix C, aggregating the SoniDo features with the shallow attention layer and token-out augmentation led to performance improvement.
@@ 4 Experiments
We conducted experiments to examine the usefulness of features extracted from music foundation models for understanding and generative downstream tasks by addressing two questions: Q1: Do extracted features have useful information for music understanding? Q2: Can extracted features boost current task-specific models for both understanding and generative tasks? To verify the generalizability of the results, we test not only the SoniDo features, but also features extracted from Jukebox and the two public versions of MusicGen (Copet et al., 2023), namely MusicGen Small and MusicGen Large. As a major focus of this work is to extend the applicable downstream tasks, we evaluate Jukebox’s features specifically for time-varying tasks and report the results from Castellon et al. (2021) for time-invariant tasks in Table 1.
To address the first question, we selected eight music tagging tasks and music transcription as representatives for understanding tasks.We verified that the extracted features encompass both time-invariant information of overall musical properties and time-varying information of specific musical events. For the second question, we tested the injection of extracted features into several task-specific models for understanding and generative tasks. They consist of one transcription model (Toyama et al., 2023), two separation models (Mitsufuji et al., 2022; Fabbro et al., 2023), and two mixing models from Martínez-Ramírez et al. (2022).
@@ 4.1 Usefulness of Extracted Features for Music Understanding
The following feature probing experiments demonstrate that features extracted from SoniDo, MusicGen and Jukebox all contain valuable information, which is consistent with Castellon et al. (2021). However, for music transcription, the shallow network remains insufficient to match SOTA models. In Section 4.2, we show that injecting extracted features into task-specific models can boost their performance beyond SOTA.
@@ 4.1.1 Music Tagging
We test a wide range of music tagging tasks as well as the emotion regression task: MagnaTagATune (MTAT) (Law et al., 2009) for auto tagging, Nsynth (Engel et al., 2017) for pitch and instrument recognition, EmoMusic (Soleymani et al., 2013) for emotion regression, GTZAN (Tzanetakis & Cook, 2002) for genre classification, GiantSteps (Knees et al., 2015; Korzeniowski & Widmer, 2017) for musical key estimation, and VocalSet (Wilkins et al., 2018) for singer and singing technique identification. A summary of the datasets is shown in Table 9 in Appendix C. We followed the pre-processing in previous studies (Li et al., 2023b; Yuan et al., 2023) and used scikit-learn (Pedregosa et al., 2011) and mir_eval (Raffel et al., 2014) for metric computation. The average R2 of arousal and valence axis is reported for EmoMusic. The feature pre-processing for time-invariant downstream tasks in Section 3.2 including the feature aggregation and token-out augmentation is applied for all tasks. Following common practice (Castellon et al., 2021; Li et al., 2023b), an MLP is then used to probe the aggregated features.
@@
Table 3: Results of feature probing using shallow back-end on MAPS (bold: best, underline: second-best).
@@ 4.1.2 Music Transcription
Beyond time-invariant understanding tasks, we continue the test on music transcription, which is a time-varying understanding task. All of SoniDo, MusicGen, and Jukebox features are obtained with unconditional extraction described in Section 3.1. The dimension and time resolution of extracted features are aligned to those of the spectrogram with linear layers. Following Castellon et al. (2021), the features are concatenated with the spectrogram. A single-layer shallow back-end network is used to probe these features.
We show the transcription performance of the feature probing mentioned above on the MAPS dataset (Emiya et al., 2010) in Table 3. All the features greatly improved the note-wise F1 score compared with using the spectrogram only. This suggests that all of SoniDo, MusicGen, and Jukebox features contain useful information for time-varying understanding tasks.
@@ 4.2 Using Extracted Features to Boost Existing Task-specific Models
In Section 4.1, we showed that the extracted features contain useful knowledge for music understanding. In this section, we test all of SoniDo, MusicGen, and Jukebox features on several SOTA task-specific models, the tasks include music transcription, music source separation, and music mixing, which covered both music understanding and generative tasks. The experimental results indicate that the extracted features consistently boost the performances of task-specific models. We also observed that injecting the SoniDo features accelerated the decrement of training loss in early epochs.
@@ 4.2.1 Music Transcription: hFT-Transformer
We applied the extracted features to hFT-Transformer (Toyama et al., 2023), a SOTA music transcription model for piano on MAPS (Emiya et al., 2010), to assess whether it surpasses existing models that rely solely on the spectrogram. On the basis of the input spectrogram, hFT-Transformer estimates the frame-based note activation, along with the onset, offset, and velocity of a note (frame, onset, offset, and velocity). It is a transformer-based model consisting of two transformer encoders that work on different axes of the input and a transformer decoder in the middle of these two encoders. Following the processing pipeline in Section 4, we attempted injecting the SoniDo, MusicGen, and Jukebox features before the 1st encoder, 2nd encoder, and decoder. We found that feature injection before the decoder yields the best result, thus we adopt this injection method in the following experiments. All the training hyperparameters were kept the same as in a previous study (Toyama et al., 2023). Further details are provided in Appendix D.3.
Table 4: Music transcription results of F1 scores on MAPS (bold: best score, underline: second best). “Note” refers to note-wise estimation. First row corresponds to hFT-Transformer (Toyama et al., 2023).
@@
Spectrogram + Jukebox 70.43 73.76 45.80 30.42
Following the common evaluation practice (Gardner et al., 2022; Toyama et al., 2023), we report four F1 scores: frame-wise, note-wise, note-wise with offset, and note-wise with offset and velocity using the checkpoint with the best validation F1 score. As shown in Table 4, injecting either SoniDo, MusicGen Small, or Jukebox features improves the performance of hFT-Transformer. The performance gap is especially huge when the model is trained with a small subset of MAPS. This demonstrates the usefulness of injecting music foundation model features into downstream task models when training data are scarce. We also observed that the decrement of training loss is faster when either SoniDo, MusicGen, or Jukebox features are injected, as shown in Figure 8 in Appendix D.3.
@@ 4.2.2 Music Source Separation: UMX, HTDemucs
We select Open-Unmix (UMX) (Stöter et al., 2019) and Demucs (HTDemucs) (Rouard et al., 2023) for feature injection. UMX estimates the time-frequency mask of the target source using recurrent neural network (RNN) blocks. HTDemucs is a hybrid model with waveform U-Net branch and spectral U-Net branches. We inject the extracted features into the encoder block for UMX using a down-sampling block and in each HTDemucs branch using a cross-domain Transformer (details in Appendices E.1 and  E.3). Based on the observation in Section 4.2.1 and for simplicity, we inject only top-level features from SoniDo.
Table 5 lists the SDR scores on the test split of MUSDB18 Rafii et al. (2017) and the hidden split of MDXDB21 Mitsufuji et al. (2022); Fabbro et al. (2023). The details of the experiments are provided in Appendix E. Similar to the experiment discussed in Section 4.2.1, a faster loss decrement is observed, as shown in Figures 12 and 13 in Appendix E. Injecting the SoniDo features into both UMX and HTDemucs greatly boosts the separation performance for both models, even on the unseen dataset MDXDB21. It also improves the separation performance of HTDemucs when training on data corrupted by bleeding errors (SDXDB23_BleedingFabbro et al. (2023). However, the MusicGen features do not always improve the results. Injecting the MusicGen Small features improved UMX, but not for the other cases. According to the ablation study results, injecting short-term Fourier transform (STFT) signal, CLAP features, MusicGen features, or Jukebox features leads to unstable training. However, no such behavior is observed when injecting the SoniDo features into HTDemucs. We assume that performance can be improved if instability during training is avoided. As mentioned in Section 4.2.1, interpreting information contained in MusicGen Large could cost too much capacity of the downstream model and result in performance degradation.